There are following options available on the market a. CALL with a strike price of 100 PLN and premium of 5 PLN b. PUT with a strike price of 100 PLN and premium of 10 PLN Is an arbitrage possible- explain your strategy? Would it be possible if you could either buy or write above options with the same characteristic-explain your strategy. [1,5 pt.] Result: NO YES
Answer #1

Using put call parity
S+P=C+K/(1+r)^t
Here, S+P=110
K/(1+r)^t<100
as K=100 and r>0 and t>0
So, S+P>C+K/(1+r)^t

Hence, we sell overvalued side
in this case lhs
and buy undervalued side
in this case rhs

Yes arbitrage is possible

Sell stock
Sell put
Buy call
Buy a bond with face value equal to 100